Introduction One of the most important tasks that chemists are faced with is the separation and identification of compounds fUnknown 1 H NMR 1 2 24 3 5 4 w 2 حا 0 PPMUnknown 1_13C NMR 60 50 40 30 20 10 PPM El-MS: 88.1 MP = -78 °C BP = 137 to 139 °CSciFinder® Unknown 1_IR 100 08 TRANSMITTANCE [%] 40 M 09 20 4000 3500 3000 1500 1000 500 2500 2000 WAVENUMBER (cm-1]Unknown 2_ H NMR 2 3 3 2 1 0 PPMUnknown 2 13C NMR 220 200 180 160 140 100 80 60 40 20 120 PPM El-MS: 86.1 MP = -39 °C BP = 102°CUnknown 2_IR SciFinder® 100 80 TRANSMITTANCE [%] 60 40 ww 20 4000 3500 3000 1500 1000 500 2500 2000 WAVENUMBER (cm-1]Unknown 3_ H NMR 2 3 3 4 2 1 0 w PPMUnknown 3_ 3C NMR 180 160 140 120 80 60 40 100 PPM 20 EI-MS: 88.1 MP = -83.6°C BP = 77.1°CUnknown 3_IR 100 com 08 TRANSMITTANCE [%] 40 09 20 4000 35000 3000 1500 1000 500 2500 2000 WAVENUMBER [cm-1]Unknown 4_ H NMR 4 1 6 1 w 2 1 0 N PPMUnknown 4_ C NMR 45 40 35 30 20 15 10 On 25 PPM El-MS: 73.1 MP = -50°C BP = 55.5°CUnknown 4_IR 100 80 TRANSMITTANCE [%] 40 09 20 4000 3500 3000 1500 1000 500 2500 2000 WAVENUMBER (cm-1]Unknown 5_ H NMR 1 2 2 3 12 10 8 6 2 0 N PPMUnknown 5 13C NMR 180 160 140 120 80 60 40 20 100 PPM El-MS: 136.1 MP = 180 to 181 °C BP = 274 to 275 °CUnknown 5_IR 100 80 TRANSMITTANCE [%] 40 more 09 prom 20 4000 3500 3000 1500 1000 500 2500 2000 WAVENUMBER (cm-1]
Answer #1

The compound 1 is Pentan-1-ol

the compound 2 is pentan-3-one

compound 3 is ethyl acetate

compound 4 is diethylamine

compound 5 is 4-methylbenzoic acid

For the analysis of the structure see the attachment

2 f HO OH H. mole culast weight = 882 b.p=138° Mip. =- 78°C # H 1H - N/M2 % 13C-NMR aus 46 ppm, broad Singlet, 14.0-H a. 63 p3 H molecular weight = 881 b. p. = 77.1 m. p. =-83.6°6 HC eltz H-NMR: 13 C-NMR a) 21 ppm b) 170 PPM C) 61 ppm 4 ppm a: 2.1 p5 botha Te moleculare Weights 1361 M.P. = 180-180°C, b. p. = 274-275°C 19 CH₃ it TH-NMRU a: 12.3 ppm, Singlet, IH, COO-H be d